OpenStudy (eric_d):

..FOLLOWING EXPANSION,FIND THE TERM STATED. (1+X)^10, 5TH TERM

OpenStudy (eric_d):

@Abmon98

OpenStudy (abmon98):

\[\left(\begin{matrix}10 \\ 0\end{matrix}\right)1^{10}+\left(\begin{matrix}10 \\ 1\end{matrix}\right)(1^{9}*x^1)+\left(\begin{matrix}10 \\ 2\end{matrix}\right)(1^8*x^2)+\left(\begin{matrix}10 \\ 3\end{matrix}\right)(1^7*x^3)+\left(\begin{matrix}10 \\ 4\end{matrix}\right)(1^6*x^4)\] \[(a+b)^n=\sum_{k}^{n}\left(\begin{matrix}n \\ k\end{matrix}\right)(a ^{n-k}*b^k)\]\[\left(\begin{matrix}n \\ k\end{matrix}\right)=n!/k!(n-k)!\] \[\left(\begin{matrix}10 \\ 0\end{matrix}\right)=10!/(10-0)!0!=1\] \[\left(\begin{matrix}10 \\ 1\end{matrix}\right)=10!/(10-1)!1!=10\] \[\left(\begin{matrix}10 \\ 2\end{matrix}\right)=10!/(10-2)!2!=45\] \[\left(\begin{matrix}10 \\ 3\end{matrix}\right)=10!/(10-3)!3!=120\] \[\left(\begin{matrix}10 \\ 4\end{matrix}\right)=10!/(10-4)!4!=210\]

OpenStudy (abmon98):

i think you can now continue till the term x^5

OpenStudy (eric_d):

But the given answer is 210x^4 @Abmon98

OpenStudy (abmon98):

\[\left(\begin{matrix}10 \\ 5\end{matrix}\right)(1^5*x^5)+\left(\begin{matrix}10 \\ 6\end{matrix}\right)(1^4*x^6)+\left(\begin{matrix}10 \\ 7\end{matrix}\right)(1^3*x^7)+\left(\begin{matrix}10 \\ 8\end{matrix}\right)(1^2*x^8)+\left(\begin{matrix}10 \\ 9\end{matrix}\right)(1*x^9)+x ^{10}\]

OpenStudy (abmon98):

Count from 1^10 5 terms

OpenStudy (abmon98):

you will find out that the answer is 210x^4

OpenStudy (eric_d):

Okay Got it Thanks @Abmon98

OpenStudy (abmon98):

your welcome :')